OpenStudy (anonymous):

Simplify the expression 2x(3+5-1)x2 A) 14 B) 21 C) 28 D)30 Please help!

OpenStudy (anonymous):

The answer is C.) 28. You do the parenthesis first. Then the multiplication is commutative.

OpenStudy (anonymous):

3+5 = 8 8 - 1 = 7 2 x 7 = 14 14 x 2 is 28.

OpenStudy (anonymous):

28

OpenStudy (anonymous):

2x(3+5-1)X2 3+5=8 2x(8-1)X2 2x(7)X2 2X7X2 14X7=28x